How much of a 35% alcohol solution shoud be mixed with 2 liters of a 60% alcohol solution to produce a 50% alcohol solution?
Answer by checkley71(8403)

.35X+.60*2=.50(X+2)
.35X+1.2=.5X+1
.35X-.5X=1-1.2
-.15X=-.2
X=-.2/-.15
X=1.3333 LITERS OF 35% ALCOHOL IS NEEDED.
PROOF
.35*1.3333+.6*2=.5(1.3333+2)
.4667+1.2=.5(3.3333)
1.667=1.667